Calculate the force between two parallel wires that are 4.7 m long and each carries a current of 2 amps. The separation between the wires is 5.0 mm.

(permeability is 4 p x10-7 Tm/A)  Answer in units of Newtons.
Answer #1

Since, the formula of magnetic field due to a long wire carrying current 'i' at a distance 'd' is that,

B = \frac{\mu_0 i}{2 \pi d}

hence magnetic field due to first wire carrying current i1 at a distance d is,

B = \frac{\mu_0 i_1}{2 \pi d} ..................................(1)

since the formula of magnetic force is,

F = iBl \sin\theta .....................................(2)

hence the force between the two parallel wires is that,

F = i_2(\frac{\mu_0 i_1}{2 \pi d})l \sin 90 \degree..............(3)

given that, i1 = i2 = 2 Amp , l = 4.7 m and d = 5 x 10-3 m.

hence force between the wires is,

F = \frac{\mu_0 i_1 i_2 l}{2 \pi d} = \frac{4 \pi \times 10^{-7} \times 2 \times 2 \times 4.7}{2 \pi \times 5 \times 10^{-3}}

F = 7.52 \times 10^{-4} Newtons..و e. مگر کی۔ - و -